The Howard family and the Gray family each used their sprinklers last summer. The water output rate for the Howard family's sprinkler was
40L per hour. The water output rate for the Gray family's sprinkler was
25L per hour. The families used their sprinklers for a combined total of
45 hours, resulting in a total water output of
1500L. How long was each sprinkler used?
Howards family sprinklers:___
Grays family sprinkler:___

Found 2 solutions by lwsshak3, macston:
Answer by lwsshak3(11628)

The Howard family and the Gray family each used their sprinklers last summer. The water output rate for the Howard family's sprinkler was
40L per hour. The water output rate for the Gray family's sprinkler was
25L per hour. The families used their sprinklers for a combined total of
45 hours, resulting in a total water output of
1500L. How long was each sprinkler used?
Howards family sprinklers: 25 hrs
Grays family sprinkler: 20 hrs
***
let x=time Howards family sprinklers were used
45-x=time Grays family sprinkler were use
..
rate*time used=output
40x+25(45-x)=1500
40x+1125-25x=1500
15x=375
x=25
45-x=20

Answer by macston(5194)

H=Howards; G=Grays
G+H=45
H=45-G
40H+25G=1500 Substitute for H
40(45-G)+25G=1500
1800-40G+25G=1500 Subtract 1800 from each side,
-15G=-300 Divide each side by -15
G=20 ANSWER 1: The Grays sprinkler was used 20 hours.
H=45-G=45-20=25 ANSWER 2: The Howards sprinkler was used 25 hours.
CHECK:
40H+25G=1500
40(25)+25(29)=1500
1000+500=1500
1500=1500